what is the standard form of y^2+4x+4y=o?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Solve for in terms of . That is, move all of the terms to the right side and then divide by the coefficient on


John


Answer by stanbon(75887)

what is the standard form of y^2+4x+4y = 0?
--------------------------------------------
Complete the square on the y-terms:
y^2 + 4y +4 = -4x+4
(y+2)^2 = -4(x-1)
======================
Cheers,
Stan H.